# Lindblad-like quantum tomography for non-Markovian quantum dynamical maps

### v1

@@ IV Conclusions and outlook
We have presented $\ell$QT, a new tool designed to characterize non-Markovian dephasing noise in QIPs. Building upon the established framework of Lindblad quantum tomography, $\ell$QT extends the applicability of Lindblad learning to scenarios where temporal correlations and non-Markovian dynamics play a significant role. In particular, it allows us to extend the characterization of the generators of quantum dynamical maps that go beyond the time-homogeneous Lindblad limit, which connect to a time-local master equation that can display negative decay rates in certain time intervals and, thus, strictly non-Markovian quantum evolutions. Through a detailed comparative study, both frequentist and Bayesian approaches to $\ell$QT are presented, offering insights into the accuracy and precision of noise estimation under different conditions.
@@
We have compared the performance of both approaches for two different dephasing quantum dynamical maps, either for a semi-classical or for a quantum-mechanical noise model. In both cases, the microscopically-motivated parametrization allows one to interpolate between a fully Markovian Lindblad limit, for which we derive analytical solutions for the optimal estimation, and a time-correlated and even non-Markovian regime which require a different distribution of the optimal and Bayesian measuring times. Interestingly, in the quantum-mechanical dephasing model, which can be obtained from a microscopic model of a qubit that is coupled to a dissipative bosonic mode in both superconducting-circuit and trapped-ion architectures, the best of the two approaches depends on whether we are in the Markovian or the non-Markovian regime. The Bayesian approach yields much better results in the non-Markovian regime, showing that it is able to automatically adapt to the particularities of the non-Markovian evolution to make much better estimations with a limited number of shots. Moreover, we also compare to more standard schemes considered in the context of Lindbladian quantum tomography, in which the measurements are distributed uniformly, showing an advantage of our schemes that again becomes more appreciable in the non-Markovian regime.
@@
Acknowledgements.
Appendix A Time-local master equation for pure dephasing
For the sake of completeness, we present here a derivation of the time-local master equation (12) for a qubit subjected to time-correlated dephasing noise, both in a semi-classical and a fully quantum-mechanical model. This serves to introduce well-known concepts and set our notation following [77].
Semi-classical time-correlated dephasing.-
The qubit evolves under a stochastic rotating-frame Hamiltonian
@@ Appendix A Time-local master equation for pure dephasing
where $\delta\tilde{\omega}(t)$ is the detuning of the qubit with respect to the frequency of a driving used in the initialization/measurement stages with respect to, and we have set $\hbar=1$. We use a tilde to highlight the random nature of $\delta\tilde{\omega}(t)$, which is modeled as a stochastic process with zero mean $\mathbb{E}[\delta\tilde{\omega}(t)]=0$, thus assuming that the driving frequency is resonant with the qubit transition on average. We recall that the averages are taken with respect to the underlying joint probability density function (pdf) of the process for any finite set of times $p_{\delta\tilde{\omega}}(\delta\bm{\omega})=p_{t_{1},t_{2},\cdots,t_{n}}(\delta\omega_{1},\delta\omega_{2},\cdots,\delta\omega_{n})$, $\forall n\in\mathbb{Z}^{+}:\delta\omega_{n}=\delta\omega(t_{n})\in\mathbb{R},\,\{t_{i}\}_{i=1}^{n}\in T$, which fulfills the conditions $p_{\delta\tilde{\omega}}(\delta\bm{\omega})\geq 0$ and $\int\prod_{n}{\rm d}\delta\omega_{n}p_{\delta\tilde{\omega}}(\delta\bm{\omega})=1$  [120, 111]. Physically, these stochastic fluctuations can either stem from frequency/phase noise of the drive, or from additional external fields that shift the energy of the qubit. For each individual trajectory of the noise $\delta\tilde{\omega}(t)$, the evolution of an initial qubit state $\rho_{0}=\ket{\psi_{0}}\!\bra{\psi_{0}}$ | in the rotating frame is purely unitary but random, giving rise to $\tilde{\rho}(t)$, and expectation values will thus depend on stochastic averages $\mathbb{E}$, leading to a completely-positive trace-preserving (CPTP) map after averaging $\rho(t)=\mathbb{E}[\tilde{\rho}(t)]=\mathcal{E}_{t,t_{0}}({\rho}_{0})$ [1, 121, 122, 64]. The corresponding stochastic differential equations are
@@
where we have defined a characteristic time as $T_{2}=2/S(0)$. The cluster property for the higher $n$-th order contributions, which have $(n-1)$ nested integrals, states that the corresponding kernels scale with $\mathcal{K}_{n}(t)\sim\alpha\zeta^{n-1}$, justifying a low-order truncation whenever the condition $\zeta\ll 1$ is met. This is known as a fast-fluctuation expansion and, back from the rotating frame, yields the time-local master equation (12).
Before finishing this section of the Appendix, let us note that the above truncation rests on the importance of the memory effects $\tau_{\rm c}$ within the $T_{2}$ time. As discussed in more detail in the main text, this $T_{2}$ time controls the time scale for the decay of coherences $\langle\sigma_{x}(t)\rangle\approx\mathord{\rm e}^{-t/T_{2}}$ in a long-time Lindbladian limit $t\gg\tau_{\rm c}$. However, for shorter times, the structure of the noise can actually lead to deviations from this limit, leading to a coherence decay that is not exponential. As emphasized in the main text, this is not an univocal signal of non-Markovianity for the qubit evolution. We note that there is an exception to the $\tau_{\rm c}\ll T_{2}$ requirement for Gaussian random processes, which are defined by a joint pdf that is a multivariate normal distribution for any set of times. In this case, the time-local master equation (12) is actually an exact result, independently of the value of $\zeta$. In fact, the higher-order contributions to the kernel [74] vanish identically $\mathcal{K}_{n}(t)=0,\forall n>2$, due to Isserlis’ theorem, most commonly referred to as Wick’s theorem in the context of physics $\mathbb{E}[\delta\tilde{\omega}(t_{1})\delta\tilde{\omega}(t_{2})\cdots\delta\tilde{\omega}(t_{n})]=\sum_{\sigma\in S_{n}}\mathbb{E}[\delta\tilde{\omega}(t_{\sigma(1)})\delta\tilde{\omega}(t_{\sigma(2)})]\cdots\mathbb{E}[\delta\tilde{\omega}(t_{\sigma(n-1)})\delta\tilde{\omega}(t_{\sigma(n)})]$, where ${S}_{n}$ is the group of all possible permutations of $n$ elements, e.g. $\sigma(1,2,\cdots,n-1,n)=(n,1,2,\cdots,n-1)$.
Quantum-mechanical time-correlated dephasing.-
We consider a single qubit coupled to an environment, and evolving ${\rm d}\rho_{\rm SB}/{\rm d}t=\mathcal{L}_{t}(\rho_{\rm SB})$ under the following Liouvillian
@@
where $B(t)$ is an environment/bath operator that introduces fluctuations on the qubit frequency, and $\mathcal{L}_{\rm B}(\bullet)$ is the Liouvillian of the bath. In the standard description of quantum master equations, the environment is macroscopically large and subject to a purely-unitary evolution $\mathcal{L}_{\rm B}(\bullet)=-\mathord{\rm i}[H_{\rm B},\bullet]$. When the system-environment coupling is weak, one can assume that the environment remains unaltered, such that the evolution takes place on the qubit but there is no back action $\rho_{\rm SB}=\rho(t)\otimes\rho_{B}$. A Born-Markov approximation then yields a non-unitary master equation for the qubit [119]. This can be expressed as a time convolutionless master equation as the one discussed in the previous subsection (54), also truncated at second order, where $\mathcal{P}$ is now a super-operator tracing over the bath degrees of freedom $\mathcal{P}(\rho_{\rm SB})=\mathrm{Tr}_{B}(\rho_{\rm SB})=\rho$  [74]. The non-unitary evolution of the qubit results from the large number of degrees of freedom in the environment, such that the purity of the state can only decrease with no recurrences.
@@
$C(t,t^{\prime})=\big{\langle}B_{I}(t)B_{I}(t^{\prime})\big{\rangle}=\mathrm{Tr}_{B}\{B_{I}(t)B_{I}(t^{\prime})\rho_{\rm B}^{\rm ss}\}.$ (63)
Here, we have assumed that $\mathcal{P}({B}_{I}(t))=0$, and we note that $B(t)$ need not commute with itself at different times. Once more, if these quantum-mechanical auto-correlation functions are wide-sense stationary, $C(t,t^{\prime})=C(t-t^{\prime})$. We note that, in contrast to the semi-classical case where $S(\omega)=S(-\omega)$, this is not necessarily the case in the quantum-mechanical case $S(\omega)\neq S(-\omega)$ [113]. However, in the case of pure dephasing, the time evolution (58) only depends on the symmetrized auto-correlation function $\bar{C}(t,t^{\prime})=\frac{1}{2}(C(t,t^{\prime})+C(t^{\prime},t))$ and therefore only the symmetric part of the auto-correlation function influences the time evolution
@@
Appendix B Analytical results for the frequentist and Bayesian estimation of Markovian Lindblad dephasing
In this appendix, we provide an analytical solution of both the frequentist and Bayesian estimation for the LQT of single-qubit Lindblad dephasing. This serves to benchmark some of the limiting results of Sec. III.3. In the Lindblad approximation to pure dephasing, we can substitute $\gamma(t)=\gamma$ in Eq. (12), obtaining a constant dephasing rate. This leads to $\Gamma(t)=t/T_{2}$ in Eq. (18), where the dephasing time is $T_{2}=1/2\gamma$, such that the likelihood function $p^{\rm L}_{i}(m_{x}|T_{2})=\frac{1}{2}(1+(-1)^{m_{x}}\mathord{\rm e}^{-t_{i}/T_{2}})$ shows a purely-exponential coherence decay. Assuming that the actual qubit dephasing follows this pdf with a real value of $T_{2\star}$, or equivalently $\gamma_{\star}$, we have a single noise parameter to learn, and could thus estimate it by measuring at a single time $t_{i}=t$. Note that the relative frequencies in this case would follow $\tilde{f}_{i}(m_{x}|T_{2\star})=\frac{1}{2}(1+(-1)^{m_{x}}\mathord{\rm e}^{-t_{i}/T_{2\star}})$ for $N_{\rm shot}\gg 1$.
[…]
We note that, in the context of Ramsey estimators for frequency standards, an atomic qubit will acquire a phase that depends on the detuning of its frequency with respect to that of the field that drives the initial and final $\pi/2$ rotations in the initialization and measurement steps [129, 130, 131, 132]. In the absence of dephasing, both the signal and the error have the same sinusoidal dependence on the elapsed time, and the signal to noise decays with $1/t_{i}\sqrt{N_{\rm shot}}$ regardless of the time at which one measures, leading to the aforementioned standard quantum limit. If there are other sources of noise beyond shot noise, this is no longer the case, and the error is minimized by maximizing the slope of the signal, i.e. maximizing the Fisher information (30).
Let us now switch to the Bayesian approach in which, at each step, we select the best time for the next measurement considering our current knowledge of the parameters. We here show that these optimal times can be obtained analytically for the Lindbladian evolution. We will assume that we are trying to determine $\gamma_{\star}=1/2T_{2\star}$ and the prior probability distribution for our knowledge about $\gamma_{\star}$ is normal $\pi_{0}(\gamma)\sim{N}(\hat{\gamma}_{0},\sigma_{0}^{2})$ with a small dispersion around the mean $\sigma_{0}\ll\hat{\gamma}_{0}$. Therefore, the parameter space is $\theta=\gamma\in\Theta=\mathbb{R}^{+}$, and the likelihood function for obtaining a single new outcome $\delta\mathbb{D}_{\ell}=\{m_{x}\}$ conditioned on our statistical knowledge of the noise is $p_{\rm TL}(\delta\mathbb{D}_{\ell}|\gamma_{\ell})=\frac{1}{2}(1+(-1)^{m_{x}}\mathord{\rm e}^{-2\gamma_{\ell}t})$. We can thus calculate the posterior $\pi_{\ell}(\gamma|\mathbb{D}_{\ell-1}\cup\delta\mathbb{D}_{\ell})$ at each $\ell>0$ step using the Bayesian update in Eq. (34) for each of the two possible outcomes $m_{x}\in\{0,1\}$. Accordingly, the expected variance of the updated distribution can be expressed as follows
[…]
Minimizing the expected variance of the posterior produced similar results to the maximization of the information gain of Eq. (37) in the case of a normal prior. Moreover, assuming that the priors also follow a Gaussian distribution $\pi_{\ell}\sim N(\hat{\gamma}_{\ell},\sigma^{2}_{\ell})$, which is justified in the asymptotic limit of a large number of shots, one can solve the integrals analytically [133], obtaining a simple update rule for the expected variances
Let us close this appendix by discussing how $\sigma_{\ell}^{2}\approx\mathbb{E}[\sigma_{\ell}^{2}]$ scales with the Bayesian step according to Eq. (68). We also assume that $2\hat{\gamma}_{\ell-1}t_{\ell}\gg\sigma^{2}_{\ell-1}t^{2}_{\ell}$, and that we have applied a sufficiently large number of Bayesian steps such that the changes in our knowledge are small and $\hat{\gamma}_{\ell}\simeq\hat{\gamma}_{\ell-1}$, all of which is reasonable in the asymptotic regime when we have done many measurements and $\sigma^{2}_{\ell-1}$ is small. We note that there might be small deviations of $\sigma_{\ell}^{2}$ from $\mathbb{E}[\sigma_{\ell}^{2}]$ but, after some additional measurements, they should eventually become vanishingly small. Under these assumptions, we have $\sigma^{2}_{\ell}=\sigma^{2}_{\ell-1}-k(t_{\ell},\hat{\gamma}_{\ell-1})\sigma_{\ell-1}^{4}$, with $k(t_{\ell},\hat{\gamma}_{\ell-1})=t_{\ell}^{2}/(\mathord{\rm e}^{2\hat{\gamma}_{\ell-1}t_{\ell}}-1)$. The recursive sequence $a_{\ell}=a_{\ell-1}-ca_{\ell-1}^{2}$ scales as $O(1/\ell)$ for $c$ a positive constant, $ca_{0}\ll 1$ and $0<a_{0}\ll 1$. Therefore $\sigma_{\ell}^{2}$ will scale as $O(1/\ell)$ independently of the time $t$ we choose to measure (i.e. we assume here that one chooses the same time $t$ for every step $\ell$). Therefore, differences between the Bayesian optimal approach and any other inference that chooses non-optimal times can only be noticed when the variance of the prior is still large. Once we have taken sufficient steps, and the standard deviation has decreased to a sufficiently small value, both the optimal measurement and the non-optimal measurement should look like two parallel lines in a log-log plot of standard deviation against $\ell$, see Fig. 11.
Figure 11: Standard deviation of the posterior $\sigma_{\ell}$ at each Bayesian step for the dephasing rate $\gamma$ in Markovian dephasing. The blue line represents the standard Bayesian approach in which at each step we find the optimal time using Eq. (37). The orange line simply chooses a random time $t_{\ell}\in\left[0.1,3\right]$. In both cases we start from the same uniformly distributed prior $U(\gamma_{\star}/3,3\gamma_{\star})$. At the beginning, the standard approach obtains an advantage, later on, once the number of measurements is sufficiently large, both instances become straight lines with slope $-\frac{1}{2}$ in this log-log plot. A total of 30 Bayesian runs were done in each case to estimate the expected $\sigma_{\ell}$. 50 measurement shots were taken at each step.
Appendix C Comparison to uniform-time Ramsey estimation
As discussed at the end of Sec. III.3.2, one can also compare the efficiency of the estimators introduced in the main text with a different one in which the measurement times are uniformly distributed over the whole time interval $T$. Such time distributions have been considered in previous works for Lindblad quantum tomography [87] and spin-locking spectroscopy of correlated noise [114]. In this appendix, we explore them in the context of the non-Markovian quantum dephasing.
This comparison can be done by calculating the determinants of the covariance matrix of the estimators, as is shown in Fig. 12 for two examples with different microscopic parameters. One parameter set shows a Markovian evolution, while the other leads to a non-Markovian one. The frequentist estimator with three optimal times is depicted by solid lines, and is always the best one in the asymptotic limit of an infinitely-large number of measurements $N_{\rm shot}\gg 1$. However, before reaching this regime, the uniform estimator displayed by dashed lines can become better. We are specially interested in the region with $10^{4}<N_{\rm shot}<10^{5}$, where the variance is already good enough to stop making more measurements, but the cost in terms of number of measurements is still not too big. In this region and for the Markovian case (blue), the uniform estimator is better than the frequentist one, which only considers three evolution times, set at the optimal values of the asymptotic limit. For the non-Markovian case (orange), the comparison is the opposite.
Figure 12: The efficiency of the optimal (solid) and uniform (dashed) estimators displayed in terms of $\det\mathrm{Cov}(\bm{\hat{\theta}}_{\rm F})^{1/6}$, with parameters $\bm{\theta}=(g^{2}_{\bar{n}},\kappa,\Delta_{\rm c})$ for different number of total measurements $N_{\rm shot}$. We consider two cases: (i) the Markovian case (blue) in which $\tau_{\rm c\star}=2/\kappa=0.3T_{2\star}$, $\Delta_{\rm c\star}=5/T_{2\star}$, (ii) the non-Markovian case (orange) where $\tau=2/\kappa=0.9T_{2\star}$, $\Delta_{\rm c}=5/T_{2\star}$. $T_{2}$ is set to 1, which makes $g^{2}\bar{n}\approx 2.7$ for the first and $g^{2}\bar{n}\approx 5.9$ for the second case. The uniform estimator measures time in 20 different equidistant points in the interval $t\in[0.02,3]$. Asymptotically $\det\mathrm{Cov}(\bm{\hat{\theta}}_{\rm F})^{1/6}$ scales as $1/\sqrt{N_{\rm shot}}$ for both estimators. For each value $N_{\rm shot}$, data is sampled and fitted to obtain an estimation. This is done a total of 5000 times to obtain afterwards the expected value of the covariance at that point. Bounds for the least squares algorithm were set to $(\bm{\theta}_{\!\star}/3,3\bm{\theta}_{\!\star})$. The initial guess was randomly chosen in that range.
Figure 13: Ratios between determinants of covariance matrices of different estimators for different parameter values $\bm{\theta}_{\!\star}=(g^{2}_{\bar{n}\star},\kappa_{\star},\Delta_{\rm c\star})$ and a fixed number of total measurements $N=2\cdot 10^{4}$. $T_{2}=1$ for all combinations of $\tau$ and $\Delta_{\rm c}$. (a) The ratio between the uniform and the optimal estimator $r_{\mathrm{uni,opt}}=(\det\mathrm{Cov}(\bm{\hat{\theta}}_{\rm F,uni})/\det\mathrm{Cov}(\bm{\hat{\theta}}_{\rm F,opt}))^{1/6}$. (b) Ratio between the uniform estimator and the Bayesian approach, $r_{\mathrm{uni,Bay}}=(\det\mathrm{Cov}(\bm{\hat{\theta}}_{\rm F,uni})/\det\Sigma_{\ell_{\max}})^{1/6}$. For the frequentist estimators the same settings as in Fig. 12 were used to determine the expected value of the determinant. For the Bayesian approach a total of 30 experiments were run, with starting prior given by the continuous uniform distribution $U(\bm{\theta}_{\!\star}/3,3\bm{\theta}_{\!\star})$, i.e., the same bounds that were given to the least squares algorithm. The dashed white line separates the Markovian from the non-Markovian regime.
In order to extend this comparison between the uniform and the optimal-time estimators for a broader set of parameters, we compute the ratios between the determinant of the covariance matrix of the two estimators, $r_{\mathrm{uni,opt}}=\det\mathrm{Cov}(\bm{\hat{\theta}}_{\rm F,uni})^{1/6}/\det\mathrm{Cov}(\bm{\hat{\theta}}_{\rm F,opt})^{1/6}$, for different values of parameters $\tau_{\rm c\star}$ and $\Delta_{\rm c\star}$, as shown in Fig. 13a and analogously to what we do in Fig. 10 of Sec. III.3.2. This determinant ratios are computed for $N_{\rm shot}=2\cdot 10^{4}$, before reaching the region with asymptotic behavior, where the optimal estimator is always better than the uniform one. We can clearly see that the uniform estimator outperforms the optimal one in the Markovian regime (upper right region), while the opposite happens in the non-Markovian regime. While the optimal estimator is not realistic, since we need to know beforehand the true value of the parameters so that we can determine the optimal times, we may use it as a baseline to compare against other estimators or try to approximate it if we have some prior knowledge of the parameters. This comparison shows that the uniform estimator might be a good choice in the non-Markovian regime when doing a small number of shots.
Appendix D Error analysis and asymptotic statistics
@@
We show here the asymptotic normality of the maximum-likelihood estimator (27) when $N_{\rm shot}$ tends to infinity and how this is related to the Fisher information in Eq. (31). The asymptotic covariance matrix is used in the main text to obtain the optimal times of the frequentist estimator and can be used to obtain approximate confidence intervals of estimations.
@@
After the linear transformation (75), we obtain that $\Delta\bm{\theta}$ behaves as

### v2

@@ Discussion
We have presented $\ell$QT, a new tool designed to characterize non-Markovian dephasing noise in QIPs, building upon the established framework of Lindblad quantum tomography. $\ell$QT extends the applicability of Lindblad learning to scenarios where temporal correlations and non-Markovian dynamics play a significant role. In particular, it allows us to extend the characterization of the generators of quantum dynamical maps that go beyond the time-homogeneous Lindblad limit, which connect to a time-local master equation that can display negative decay rates in certain time intervals and, thus, strictly non-Markovian quantum evolutions. Through a detailed comparative study, both frequentist and Bayesian approaches to $\ell$QT are presented, offering insights into the accuracy and precision of noise estimation under different conditions.
@@
We have compared the performance of both approaches for two different dephasing quantum dynamical maps, either for a semi-classical or for a quantum-mechanical noise model. In both cases, the microscopically-motivated parametrization allows one to interpolate between a fully Markovian Lindblad limit, for which we derive analytical solutions for the optimal estimation, and a time-correlated and even non-Markovian regime which require a different distribution of the optimal and Bayesian measuring times. Interestingly, in the quantum-mechanical dephasing model, which can be obtained from a microscopic model of a qubit that is coupled to a dissipative bosonic mode in both superconducting-circuit and trapped-ion architectures, the best of the two approaches depends on whether we are in the Markovian or the non-Markovian regime. The Bayesian approach yields much better results in the non-Markovian regime, showing that it is able to automatically adapt to the particularities of the non-Markovian evolution to make much better estimations with a limited number of shots. Moreover, we also compare to more standard schemes considered in the context of Lindbladian quantum tomography, in which the measurements are distributed uniformly (see Sec. II of the Supplementary Material). In this case we show an advantage of our schemes that again becomes more appreciable in the non-Markovian regime for the Bayesian approach. However, we also find that for the frequentist approach, the optimal distribution of measurements does not always outperform the uniform distribution when the number of measurments is small and the asymptotic regime has not yet been reached. These results may suggest that the observed patterns extend to other similar cases, but further exploration would be required to confirm this.
The semi-classical dephasing model we have discussed has also been considered in the context of central spin models, where the central spin/qubit is dipolarly-coupled to an ensemble of environmental spins that play the role of a mesoscopic bath. For the case of NV centers in diamond [130], the bath corresponds to nearby substitutional nitrogen atoms, also known as P1 centers, which have a much smaller splitting than the qubit. Due to the mismatch, energy exchange processes can be neglected, and only longitudinal ones survive leading to a time-correlated dephasing. For these systems, assuming incoherent dynamics of different bath spins and negligible back action is a good approximation, such that the above semi-classical model with an Ornstein-Uhlenbeck process turns out to be a reasonably-good approximation. More generally, if the bath consists of weakly-interacting nuclear spins that get dipolarly-coupled to the qubit, bath correlations might be relevant. Although this is not generally valid, when the leading bath correlations can be reduced to pairs of spins, a similar Gaussian semi-classical model can still be used [131], albeit considering other PSDs. In more general situations, one should consider larger groups of spins, and explore other spin-bath dephasing models. Although this goes beyond the scope of this work, we believe that incorporating a detailed microscopically-motivated parametrization of this type of non-Markovian dephasing should allow for a more efficient parametrization of the noise that could be incorporated in a $\ell$QT that is similar in spirit to the one considered in this work.
@@
Methods
Time-local master equation for pure dephasing
For the sake of completeness, we present here a derivation of the time-local master equation in Eq. (20) for a qubit subjected to time-correlated dephasing noise, both in a semi-classical and a fully quantum-mechanical model. This serves to introduce well-known concepts and set our notation following [77].
a. Semi-classical time-correlated dephasing.- The qubit evolves under a stochastic rotating-frame Hamiltonian
@@ Methods
where $\delta\tilde{\omega}(t)$ is the detuning of the qubit with respect to the frequency of a driving used in the initialization/measurement stages with respect to, and we have set $\hbar=1$. We use a tilde to highlight the random nature of $\delta\tilde{\omega}(t)$, which is modeled as a stochastic process with zero mean $\mathbb{E}[\delta\tilde{\omega}(t)]=0$, thus assuming that the driving frequency is resonant with the qubit transition on average. We recall that the averages are taken with respect to the underlying joint PDF of the process for any finite set of times $p_{\delta\tilde{\omega}}(\delta\bm{\omega})=p_{t_{1},t_{2},\cdots,t_{n}}(\delta\omega_{1},\delta\omega_{2},\cdots,\delta\omega_{n})$, $\forall n\in\mathbb{Z}^{+}:\delta\omega_{n}=\delta\omega(t_{n})\in\mathbb{R},\,\{t_{i}\}_{i=1}^{n}\in T$, which fulfills the conditions $p_{\delta\tilde{\omega}}(\delta\bm{\omega})\geq 0$ and $\int\prod_{n}{\rm d}\delta\omega_{n}p_{\delta\tilde{\omega}}(\delta\bm{\omega})=1$  [132, 121]. Physically, these stochastic fluctuations can either stem from frequency/phase noise of the drive, or from additional external fields that shift the energy of the qubit. For each individual trajectory of the noise $\delta\tilde{\omega}(t)$, the evolution of an initial qubit state $\rho_{0}=\ket{\psi_{0}}\!\bra{\psi_{0}}$ | in the rotating frame is purely unitary but random, giving rise to $\tilde{\rho}(t)$, and expectation values will thus depend on stochastic averages $\mathbb{E}$, leading to a completely-positive trace-preserving (CPTP) map after averaging $\rho(t)=\mathbb{E}[\tilde{\rho}(t)]=\mathcal{E}_{t,t_{0}}({\rho}_{0})$ [1, 133, 134, 64]. The corresponding stochastic differential equations are
@@
where we have defined a characteristic time as $T_{2}=2/S(0)$. The cluster property for the higher $n$-th order contributions, which have $(n-1)$ nested integrals, states that the corresponding kernels scale with $\mathcal{K}_{n}(t)\sim\alpha\zeta^{n-1}$, justifying a low-order truncation whenever the condition $\zeta\ll 1$ is met. This is known as a fast-fluctuation expansion and, back from the rotating frame, yields the time-local master equation in Eq. (20).
Let us note that the above truncation rests on the importance of the memory effects $\tau_{\rm c}$ within the $T_{2}$ time. As discussed in more detail in the “Results”, this $T_{2}$ time controls the time scale for the decay of coherences $\langle\sigma_{x}(t)\rangle\approx\mathord{\rm e}^{-t/T_{2}}$ in a long-time Lindbladian limit $t\gg\tau_{\rm c}$. However, for shorter times, the structure of the noise can actually lead to deviations from this limit, leading to a coherence decay that is not exponential. As emphasized in the main text, this is not an univocal signal of non-Markovianity for the qubit evolution. We note that there is an exception to the $\tau_{\rm c}\ll T_{2}$ requirement for Gaussian random processes, which are defined by a joint PDF that is a multivariate normal distribution for any set of times. In this case, the time-local master equation in Eq. (20) is actually an exact result, independently of the value of $\zeta$. In fact, the higher-order contributions to the kernel [74] vanish identically $\mathcal{K}_{n}(t)=0,\forall n>2$, due to Isserlis’ theorem, most commonly referred to as Wick’s theorem in the context of physics $\mathbb{E}[\delta\tilde{\omega}(t_{1})\delta\tilde{\omega}(t_{2})\cdots\delta\tilde{\omega}(t_{n})]=\sum_{\sigma\in S_{n}}\mathbb{E}[\delta\tilde{\omega}(t_{\sigma(1)})\delta\tilde{\omega}(t_{\sigma(2)})]\cdots\mathbb{E}[\delta\tilde{\omega}(t_{\sigma(n-1)})\delta\tilde{\omega}(t_{\sigma(n)})]$, where ${S}_{n}$ is the group of all possible permutations of $n$ elements, e.g. $\sigma(1,2,\cdots,n-1,n)=(n,1,2,\cdots,n-1)$.
b. Quantum-mechanical time-correlated dephasing.- We consider a single qubit coupled to an environment, and evolving ${\rm d}\rho_{\rm SB}/{\rm d}t=\mathcal{L}_{t}(\rho_{\rm SB})$ under the following Liouvillian
@@
where $B(t)$ is an environment/bath operator that introduces fluctuations on the qubit frequency, and $\mathcal{L}_{\rm B}(\bullet)$ is the Liouvillian of the bath. In the standard description of quantum master equations, the environment is macroscopically large and subject to a purely-unitary evolution $\mathcal{L}_{\rm B}(\bullet)=-\mathord{\rm i}[H_{\rm B},\bullet]$. When the system-environment coupling is weak, one can assume that the environment remains unaltered, such that the evolution takes place on the qubit but there is no back action $\rho_{\rm SB}=\rho(t)\otimes\rho_{B}$. A Born-Markov approximation then yields a non-unitary master equation for the qubit [129]. This can be expressed as a time convolutionless master equation as the one discussed in the previous subsection in Eq. (51), also truncated at second order, where $\mathcal{P}$ is now a super-operator tracing over the bath degrees of freedom $\mathcal{P}(\rho_{\rm SB})=\mathrm{Tr}_{B}(\rho_{\rm SB})=\rho$  [74]. The non-unitary evolution of the qubit results from the large number of degrees of freedom in the environment, such that the purity of the state can only decrease with no recurrences.
@@
$C(t,t^{\prime})=\mathbb{E}[B_{I}(t)B_{I}(t^{\prime})]=\mathrm{Tr}_{B}\{B_{I}(t)B_{I}(t^{\prime})\rho_{\rm B}^{\rm ss}\}.$ (60)
Here, we have assumed that $\mathcal{P}({B}_{I}(t))=0$, and we note that $B(t)$ need not commute with itself at different times. Once more, if these quantum-mechanical auto-correlation functions are wide-sense stationary, $C(t,t^{\prime})=C(t-t^{\prime})$. We note that, in contrast to the semi-classical case where $S(\omega)=S(-\omega)$, this is not necessarily the case in the quantum-mechanical case $S(\omega)\neq S(-\omega)$ [123]. However, in the case of pure dephasing, the time evolution in Eq. (55) only depends on the symmetrized auto-correlation function $\bar{C}(t,t^{\prime})=\frac{1}{2}(C(t,t^{\prime})+C(t^{\prime},t))$ and therefore only the symmetric part of the auto-correlation function influences the time evolution
@@
Error analysis and asymptotic statistics
@@
We show here the asymptotic normality of the maximum-likelihood estimator in Eq. (9) when $N_{\rm shot}$ tends to infinity and how this is related to the Fisher information in Eq. (13). The asymptotic covariance matrix is used in the main text to obtain the optimal times of the frequentist estimator and can be used to obtain approximate confidence intervals of estimations. To simplify the notation and reduce complexity, we consider the pure dephasing case here, where the cost function is simplified. However, the conclusions remain the same, and the general case can be recovered by substituting $i\rightarrow i,s,b$ and $m_{x}\rightarrow m_{b}$. In this case, instead of considering only measurement points in time, we account for measurement configurations, including evolution times, initial states, measurement bases, and measurement outcomes.
@@
After the linear transformation of Eq. (68), we obtain that $\Delta\bm{\theta}$ behaves as